# FOCUS: Fused Observation of Channels for Unveiling Spectra

### v1

@@ Abstract
Hyperspectral imaging (HSI) captures hundreds of narrow, contiguous wavelength bands, making it a powerful tool in biology, agriculture, and environmental monitoring. However, interpreting Vision Transformers (ViTs) in this setting remains largely unexplored due to two key challenges: (1) existing saliency methods struggle to capture meaningful spectral cues, often collapsing attention onto the class token, and (2) full-spectrum ViTs are computationally prohibitive for interpretability, given the high-dimensional nature of HSI data. We present FOCUS, the first framework that enables reliable and efficient spatial–spectral interpretability for frozen ViTs. FOCUS introduces two core components: class-specific spectral prompts that guide attention toward semantically meaningful wavelength groups, and a learnable [SINK] token trained with an attraction loss to absorb noisy or redundant attention. Together, these designs make it possible to generate stable and interpretable 3D saliency maps and spectral importance curves in a single forward pass—without any gradient backpropagation or backbone modification. FOCUS improves band-level IoU by 15%, reduces attention collapse by over 40%, and produces saliency results that align closely with expert annotations. With less than 1% parameter overhead, our method makes high-resolution ViT interpretability practical for real-world hyperspectral applications—bridging a long-standing gap between black-box modeling and trustworthy HSI decision-making.
1 Introduction
[…]
@@ 1 Introduction
Hyperspectral imaging (HSI) captures dense spectral signatures across hundreds of narrow, contiguous bands, enabling fine-grained analysis of physiological traits such as stress, disease, or pigment content. This spectral richness holds enormous promise for applications in biology, agriculture, and environmental monitoring. Recently, Vision Transformers (ViTs) have shown strong performance in hyperspectral recognition tasks, due to their ability to model long-range spatial–spectral dependencies.
@@
However, such progress has not translated to interpretability. For real-world deployment, HSI models must answer not only “where” decisions are made (spatial saliency), but also “which wavelengths” are responsible—a need that current ViT interpretability tools fail to meet. Existing methods like Grad-CAM [32], LayerCAM [22], or Prompt-CAM [8] were designed for 3-channel RGB inputs, and cannot scale to the high-dimensional spectral input of HSI. They either ignore the spectral axis entirely or yield coarse spatial maps without wavelength reasoning—rendering them biologically uninformative. The core barrier is computational: unlike RGB, hyperspectral cubes often contain 200–300 channels, leading to input token sequences that are orders of magnitude larger. Since ViT attention scales quadratically with token count, spectral interpretability becomes intractable in both memory and runtime cost. In practice, this forces existing methods to avoid the full spectrum altogether—leaving a fundamental gap in saliency modeling. This is further exacerbated by the attention sink phenomenon [40], where attention collapses into dominant tokens such as [CLS]. In HSI, this issue worsens with spectral scale: the more spectral tokens, the more diluted and unstable the attention becomes. When ViTs fail to lock onto meaningful wavelength patterns, they default to class tokens, leading to spatial drift and spectral blindness (Fig. 1). For example, in plant disease detection, key reflectance signals often lie in narrow bands like the red-edge (700–750 nm). Existing methods fail to highlight such bands, producing diffuse or misleading saliency maps that misalign with expert knowledge and limit downstream utility.
@@
We address these challenges by introducing FOCUS, the first framework to provide efficient and stable spatial–spectral interpretability for frozen ViTs on hyperspectral data. FOCUS introduces two lightweight but powerful innovations. First, we design class-specific spectral prompts that encode physiologically meaningful wavelength groups (e.g., VIS, red-edge, NIR), anchoring attention in the spectral domain. Second, we introduce a learnable [SINK] token trained with an attraction loss that explicitly absorbs redundant, low-relevance attention—transforming the attention sink from a failure mode into a controlled filtering mechanism.
@@
With these components, FOCUS enables reliable interpretation of ViT-based HSI models with minimal overhead. It produces (i) a high-resolution 3D saliency cube across spatial and spectral dimensions, (ii) interpretable spectral attribution curves, and (iii) stable attention pathways without requiring test-time gradients or model modification. Extensive evaluation on two HSI benchmarks—including the new HyperLeaf2024 dataset—shows that FOCUS improves band-level IoU by 15%, reduces attention collapse by 40%, and generates domain-aligned explanations. These gains are achieved with under 1% additional parameters.
@@
  • We are the first to identify and resolve a fundamental scalability barrier for ViT-based interpretability in hyperspectral imaging (HSI). Existing methods cannot scale to full-spectrum inputs due to the exponential growth in attention computation with increasing spectral channels. Our framework offers the first practical solution for interpretable ViTs under high-dimensional spectral conditions.
@@
  • We introduce FOCUS, the first interpretability framework that jointly addresses spectral relevance and attention stability. By combining class-specific spectral prompts and a learnable [SINK] token with an attraction loss, FOCUS enables precise wavelength attribution while effectively filtering noisy or redundant attention—substantially mitigating the attention collapse phenomenon.
@@
  • FOCUS is fully gradient-free, architecture-agnostic, and compatible with frozen ViTs. It generates high-resolution spatial–spectral explanations through a single forward pass, achieving state-of-the-art interpretability with less than 1% parameter overhead—making it practical for real-world deployment in biology and agriculture.
2 Related Work
2.1 Saliency Maps for Transformers
Initial CAM-based explanations emerged from gradient-centric methods, such as Grad-CAM [32], Eigen-CAM [27], Score-CAM [37], and XGrad-CAM [16]. As Transformers became prevalent, their noisy and layer-spanning gradient flows led to the development of attention-based alternatives, including AttnRollout [1], TokenGrad [46], and LRP-ViT [4], which propagate attention weights rather than backpropagate gradients. Smooth-CAM [28] further enhances robustness via integrated smoothing. More recent works [41, 42], such as Prompt-CAM [8], introduce learnable prompts to guide explanations, while DiffCAM [24] eliminates gradient reliance by comparing feature differences with reference distributions. Yet, despite their architectural innovations, all these methods operate solely in the spatial domain, lacking any mechanism to capture spectral relevance. This omission limits their applicability in domains like HSI, where spectral reasoning is foundational.
2.2 Hyperspectral Recognition and Explainability
Early approaches to HSI classification relied on 3D CNNs [25] or combinations of 2D spatial and 1D spectral convolutions [26], aiming to jointly encode local and spectral context. Recent methods leverage Transformer architectures—SST [7], HSViT [43], SpectralFormer [19], and HyperViT [31]—to capture long-range dependencies and achieve higher accuracy. However, these models typically expose only coarse spatial attention maps and offer limited interpretability in the spectral dimension. To address this, channel-level attribution methods such as SANet [13], SSA-CAM [33], and Hyperspectral Band [45] identify influential wavelengths but discard spatial granularity. Meanwhile, prompt-efficient tuning methods like HyperPrompt [18] and SPTrack-Prompt [17] improve transferability but lack explanatory depth. To date, no approach provides joint spatial–spectral saliency for frozen HSI ViTs—leaving a critical gap in understanding model decisions from both perspectives, largely due to the prohibitive computational cost of modeling dense spectral tokens in attention-based architectures.
2.3 Attention Sink in Vision Transformers
Vision Transformers (ViTs) frequently suffer from attention sink—a collapse of the attention distribution where a handful of tokens, most notably the class token, monopolise the receptive field, degrading robustness and rendering attribution maps unstable. Existing counter‑measures fall into two principal camps: Auxiliary‑token methods, such as SpecFormer, which flattens the attention spectrum by penalising large singular values during training [21]; Register‑ViT, which attaches register tokens that compete with the class token and absorb redundant activation [9]; and EDIT, which adds editing tokens encoding fine‑grained sub‑class structure that act as controllable sinks [15]; and Pruning and rescaling approaches, including TokenPruner, which progressively excises low‑saliency tokens at inference time [38, 39], Adaptive Token Sampler (ATS), which learns an importance mask that discards uninformative heads on‑the‑fly [14], and LayerScale, which mitigates early‑layer domination by rescaling residual branches with learnable coefficients [35]. Although effective, these techniques typically require full retraining and do not provide a quantitative gauge of collapse mitigation. By appending a single [SINK] token and a lightweight attraction loss that siphons non‑discriminative attention into an explicit, interpretable reservoir, our method serves as a drop‑in module for frozen backbones and, for the first time, enables direct measurement of collapse mitigation through two metrics: Sink‑Rate—the proportion of attention mass absorbed by [SINK]—and Sink‑Consistency—the inter‑layer stability of saliency maps.
3 Method
3.1 Overview
@@
Given a hyperspectral image cube $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ with $C$ spectral bands and spatial resolution $H\times W$, our goal is twofold: (i) accurately predict a target outcome, such as plant genotype, disease status, or physiological traits, and (ii) generate an interpretable joint spatial–spectral saliency map, clearly indicating both where and which wavelengths influence model decisions.
To achieve these objectives, we extend a frozen Vision Transformer (ViT) [12] backbone by introducing three novel and intuitive modules:
@@
    Class-specific spectral prompts, explicitly guiding wavelength-aware attention (Section 3.2);
@@
    A learnable [SINK] token with an associated attraction loss to convert attention collapse [40] into a controllable noise-filtering mechanism (Section 3.3);
@@
    An attention aggregation routine, forming a unified three-dimensional saliency cube from prompt-to-patch attention (Section 3.4).
@@
Figure 2: Overview of FOCUS. A hyperspectral image is embedded and processed by a frozen ViT. FOCUS injects class-specific spectral prompts, each corresponding to meaningful spectral groups (e.g., VIS, red-edge, NIR), and introduces a dedicated [SINK] token to absorb redundant attention. Saliency is computed via prompt-to-patch attention, yielding a 3D attribution cube $\mathbf{T}\in\mathbb{R}^{H\times W\times C}$ for spatial–spectral interpretation.
3.2 Class-Specific Spectral Prompts
We partition the $C$ spectral bands into $G$ non-overlapping groups $\{\mathcal{B}_{g}\}_{g=1}^{G}$, guided by domain knowledge [20], such as visible (VIS: 400–700 nm), red-edge (700–750 nm), NIR (750–1100 nm), and SWIR (1100 nm). Let $C_{g}=|\mathcal{B}_{g}|$ denote the number of bands in group $g$.
For each class $c\in\{1,\dots,K\}$ and spectral group $g$, we introduce a learnable spectral prompt token $\mathbf{p}_{c,g}\in\mathbb{R}^{d}$. The input sequence to the first ViT layer is:
@@ 3.2 Class-Specific Spectral Prompts
where $N=H\cdot W$ is the number of spatial patches, and each spectral band is embedded using a depth-wise adapter $\boldsymbol{\Phi}\in\mathbb{R}^{C\times d}$ before patchification.
Each spectral prompt token $\mathbf{p}_{c,g}$ serves as a “spectral anchor,” exclusively interacting with patches from its assigned spectral group $\mathcal{B}_{g}$. This design ensures sharper attention and enhanced spectral interpretability compared to generic prompts [8, 18].
3.3 Sink-Aware Attention Routing
At transformer layer $\ell$ and head $h$, attention is computed via standard scaled dot-product:
@@ 3.3 Sink-Aware Attention Routing
where $\mathbf{Q},\mathbf{K}\in\mathbb{R}^{(1+G+N)\times d}$, and the [SINK] token index is set as $k_{\mathrm{sink}}=1$.
We identify a subset of low-informativeness heads $\mathcal{H}_{\text{aux}}$ [34, 10] to serve as noise absorbers. We introduce a lightweight attraction loss encouraging these heads to route maximal attention to the [SINK] token:
@@
where $\lambda$ is a small scalar (e.g., $10^{-3}$). This strategy mitigates attention collapse, enabling discriminative heads to better capture spatial–spectral signals.
3.4 3-D Prompt-CAM Generation
After the final transformer layer $L$, we aggregate prompt-to-patch attention from discriminative heads =$\mathcal{H}_{\star}=\{1,\dots,H\}\setminus\mathcal{H}_{\text{aux}}$ into a spatial–spectral saliency cube:
@@
where $g(\lambda)$ denotes the spectral group index of band $\lambda$. The resulting cube $\mathbf{T}\in\mathbb{R}^{H\times W\times C}$ simultaneously encodes spatial and spectral importance.
To facilitate intuitive interpretation, we define a spatial heatmap $M_{x,y}$ and a spectral importance curve $B_{\lambda}$ as:
@@
Both visualizations are obtained efficiently in a single forward pass, without gradients.
3.5 Training Objective and Complexity
Our complete training objective combines task-specific loss $\mathcal{L}_{\text{task}}$ (cross-entropy or MSE) and the attention attraction loss $\mathcal{L}_{\text{sink}}$:
@@ 3.5 Training Objective and Complexity
During training, the ViT backbone remains frozen, while only (i) the spectral prompts ($K\times G\times d$), (ii) the [SINK] token ($d$), and (iii) the depth-wise spectral adapter ($C\times d$) are optimized. Collectively, these parameters add fewer than 1% overhead relative to the original ViT-B/16 model (0.8M vs. 86M parameters), introducing only a marginal computational cost (approximately 2.1% additional FLOPs).
Our method requires no gradient computation during inference. A single forward pass simultaneously provides accurate predictions, a comprehensive 3D saliency cube, a clear spatial heatmap, and a precise spectral relevance curve, directly supporting downstream analysis and biological interpretation.
4 Experiments
4.1 Experimental Setup
@@
Baselines
We compare FOCUS against a comprehensive suite of 18 state-of-the-art explainability methods spanning four major categories. Classical CAM-based techniques include Grad-CAM [32], Grad-CAM++ [3], Score-CAM [37], Eigen-CAM [27], XGrad-CAM [16], Layer-CAM [22], and Ablation-CAM [11], all of which rely on gradient-based activation visualization. Transformer-specific interpretability methods include AttnRollout [1], TokenGrad [46], LRP-ViT [4], Prompt-CAM [8], and DiffCAM [24], which exploit attention flow and token saliency for ViT explanations. Prototype- and data-driven models such as ProtoPNet [6], ProtoPFormer [44], INTR [29], and RISE [30] provide example-based or perturbation-based interpretability. Finally, we also evaluate two hyperspectral-specific approaches—SANet [13] and SSA-CAM [33]—that directly model spectral importance but typically lack spatial resolution. For fair comparison, all baselines are re-implemented using a common frozen ViT-B/16 (DINOv2) backbone and applied to hyperspectral cubes preprocessed under the same pipeline. Evaluation protocols follow consistent saliency metrics across spatial, spectral, and joint dimensions, ensuring a level playing field.
Implementation Details
We freeze the pretrained ViT-B/16 backbone (DINOv2) during all experiments, training only the proposed spectral prompts, the [SINK] token, and a lightweight $1\times 1$ spectral adapter. We employ the AdamW optimizer with parameters ($\beta_{1}=0.9$, $\beta_{2}=0.95$), learning rate $5\times 10^{-4}$, and weight decay $10^{-2}$, training for a total of 100 epochs. The number of spectral prompt groups ($G$) is set to 10 based on ablation results, and the sink attraction loss weight ($\lambda$) is set to $10^{-3}$.
@@
4.2 Main Results
Table 1 presents the main quantitative results of FOCUS and competitive baselines across two challenging hyperspectral benchmarks: HyperLeaf 2024 and Tomato BLS. For each dataset, we report the most informative and widely-used metrics: classification accuracy (Acc), band-level interpretability (BIO@5), spatial localization quality (IoU), and overall explanation precision (AUPRC). All methods share an identical ViT-B/16 (DINOv2) backbone and consistent training/evaluation protocols. On both datasets, FOCUS achieves the best performance in all reported metrics, demonstrating not only strong predictive ability, but also superior spatial and spectral interpretability. Notably, on HyperLeaf 2024, FOCUS improves BIO@5 and IoU by over 15% compared to the best transformer and HSI-specific baselines. On Tomato BLS, our approach yields sharper lesion localization and more accurate band attribution, further highlighting its generalizability and robustness.
Table 1: Main results across multiple hyperspectral benchmarks. For each dataset, we report classification accuracy (Acc), band-level interpretability (BIO@5), spatial localization (IoU), and explanation precision (AUPRC). Best results are in bold. FOCUS consistently leads on both predictive and interpretability metrics.
Method HyperLeaf 2024 Tomato BLS
Acc $\uparrow$ BIO@5 $\uparrow$ IoU $\uparrow$ AUPRC $\uparrow$ Acc $\uparrow$ BIO@5 $\uparrow$ IoU $\uparrow$ AUPRC $\uparrow$
Grad-CAM [32][ICCV’17] 82.1 0.33 0.41 0.53 79.3 0.28 0.39 0.51
SANet [13][TIP’20] 83.0 0.36 0.43 0.56 80.5 0.29 0.41 0.52
Prompt-CAM [8][CVPR’25] 84.5 0.38 0.46 0.59 81.2 0.34 0.43 0.58
DiffCAM [24][CVPR’25] 85.3 0.41 0.49 0.62 82.4 0.36 0.47 0.61
FOCUS (Ours) 87.0 0.56 0.58 0.72 84.6 0.47 0.54 0.70
Table 2: Efficiency and inference overhead. FOCUS offers minimal overhead and needs no gradients.
Method Params FLOPs Grad Latency
(%) (%) (ms)

### v2

@@ Abstract
Hyperspectral imaging (HSI) provides critical details for biology and agriculture, yet interpreting Vision Transformers (ViTs) on such high-dimensional data remains a challenge. Existing methods fail for two reasons: the computational cost of processing hundreds of bands is prohibitive, and models often suffer from “attention collapse,” where focus shifts to irrelevant tokens rather than meaningful spectral signals. We propose FOCUS, a framework that enables efficient spatial-spectral interpretability for frozen ViTs. Our method introduces two lightweight components: class-specific spectral prompts that anchor attention to relevant wavelength groups, and a learnable [SINK] token that actively filters out noisy attention. Unlike prior approaches, FOCUS generates 3D saliency maps in a single forward pass without requiring gradients or retraining the backbone. Experiments show that our method improves band-level IoU by 15% and reduces attention collapse by 40%. With less than 1% parameter overhead, FOCUS makes interpretable HSI practical for real-world applications.
I Introduction
@@
Hyperspectral imaging (HSI) captures dense spectral signatures across hundreds of narrow, contiguous bands. This capability enables fine-grained analysis of physiological traits, such as stress, disease, and pigment content. Consequently, HSI has become a valuable tool in biology, agriculture, and environmental monitoring. Recently, Vision Transformers (ViTs) have achieved strong performance in hyperspectral recognition tasks by modeling long-range spatial and spectral dependencies.
@@
However, this predictive power has not translated into interpretability. For real-world deployment, HSI models must explain not only “where” decisions are made (spatial saliency) but also “which wavelengths” are responsible. Current ViT interpretability tools fail to meet this requirement. Existing methods, such as Grad-CAM [1], LayerCAM [2], and Prompt-CAM [3], are designed for 3-channel RGB inputs and do not scale to high-dimensional hyperspectral data. They typically ignore the spectral axis or produce coarse spatial maps without specific wavelength information, making them biologically uninformative. The main barrier is computational. Unlike RGB images, hyperspectral cubes often contain 200 to 300 channels, creating input token sequences that are orders of magnitude larger. Since ViT attention scales quadratically with the number of tokens, calculating full spectral interpretability is often intractable due to high memory and runtime costs [4]. As a result, existing methods usually avoid the full spectrum, which leaves a fundamental gap in saliency modeling.
This challenge is further complicated by the attention sink phenomenon [5]. In this phenomenon, attention weights collapse onto dominant tokens, such as the [CLS] token. This issue becomes more severe in HSI as the spectral scale increases. With more spectral tokens, the attention becomes more diluted and unstable. When ViTs fail to capture meaningful wavelength patterns, they tend to revert to class tokens, causing spatial drift and spectral blindness (Fig. 1). For instance, in plant disease detection, key reflectance signals often reside in narrow bands like the red-edge (700–750 nm). Existing methods fail to highlight these bands and produce diffuse saliency maps that do not align with expert knowledge, limiting their practical utility.
Figure 1: Motivation of FOCUS. Black $\times$ and green denote two classes of leaf cubes in a simplified spatial–spectral feature plane.
@@
We address these challenges with FOCUS, the first framework to provide efficient and stable spatial–spectral interpretability for frozen ViTs on hyperspectral data. FOCUS introduces two lightweight innovations. First, we design class-specific spectral prompts. These prompts encode physiologically meaningful wavelength groups, such as VIS, red-edge, and NIR, to anchor attention in the spectral domain. Second, we introduce a learnable [SINK] token trained with an attraction loss. This token explicitly absorbs redundant or low-relevance attention, which turns the attention sink from a failure mode into a controlled filtering mechanism.
@@
With these components, FOCUS enables reliable interpretation of ViT-based HSI models with minimal overhead. It produces a high-resolution 3D saliency cube across spatial and spectral dimensions, interpretable spectral attribution curves, and stable attention pathways. Importantly, it achieves this without requiring test-time gradients or backbone modifications. Extensive evaluations on two HSI benchmarks, including the new HyperLeaf2024 dataset, show that FOCUS improves band-level IoU by 15%, reduces attention collapse by 40%, and generates domain-aligned explanations. These gains are achieved with less than 1% additional parameters.
@@
  • We identify and resolve a fundamental scalability barrier for ViT-based interpretability in hyperspectral imaging. Existing methods cannot scale to full-spectrum inputs because attention computation grows exponentially with the number of spectral channels. Our framework offers the first practical solution for interpretable ViTs under high-dimensional spectral conditions.
@@
  • We introduce FOCUS, an interpretability framework that addresses both spectral relevance and attention stability. By combining class-specific spectral prompts and a learnable [SINK] token with an attraction loss, FOCUS enables precise wavelength attribution. It also effectively filters noisy or redundant attention, which substantially mitigates the attention collapse phenomenon.
@@
  • FOCUS is fully gradient-free, architecture-agnostic, and compatible with frozen ViTs. It generates high-resolution spatial–spectral explanations through a single forward pass and achieves state-of-the-art interpretability with less than 1% parameter overhead. This makes it practical for real-world deployment in biology and agriculture.
II Method
II-A Overview
@@
Given a hyperspectral image cube $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ with $C$ spectral bands and spatial resolution $H\times W$, our goal is twofold: (i) accurately predict a target outcome, such as plant genotype, disease status, or physiological traits, and (ii) generate an interpretable joint spatial–spectral saliency map, clearly indicating both where and which wavelengths influence model decisions. To achieve these objectives, we extend a frozen Vision Transformer (ViT) [6] backbone by introducing three novel and intuitive modules:
@@
    Class-specific spectral prompts, explicitly guiding wavelength-aware attention (Section II-B);
@@
    A learnable [SINK] token with an associated attraction loss to convert attention collapse [5] into a controllable noise-filtering mechanism (Section II-C);
@@
    An attention aggregation routine, forming a unified three-dimensional saliency cube from prompt-to-patch attention (Section II-D).
@@
Figure 2: Overview of FOCUS. A hyperspectral image is embedded by a frozen ViT backbone. FOCUS augments it with (i) class-specific spectral prompts that anchor meaningful band groups (e.g., VIS, red-edge, NIR) and (ii) a dedicated [SINK] token to absorb redundant attention. Prompt-to-patch attention is aggregated into a 3D saliency cube $\mathbf{T}\in\mathbb{R}^{H\times W\times C}$, enabling joint spatial–spectral interpretation. We emphasize that the comparison with Prompt-CAM is only for contrast; the figure itself focuses on the FOCUS framework.
II-B Class-Specific Spectral Prompts
We partition the $C$ spectral bands into $G$ non-overlapping groups $\{\mathcal{B}_{g}\}_{g=1}^{G}$, guided by domain knowledge [7], such as visible (VIS: 400–700 nm), red-edge (700–750 nm), NIR (750–1100 nm), and SWIR (>>1100 nm). Let $C_{g}=|\mathcal{B}_{g}|$ denote the number of bands in group $g$.
For each class $c\in\{1,\dots,K\}$ and spectral group $g$, we introduce a learnable spectral prompt token $\mathbf{p}_{c,g}\in\mathbb{R}^{d}$. The initial input sequence is:
@@ II-B Class-Specific Spectral Prompts
where $N=H\cdot W$ is the number of spatial patches. Each band of $\mathbf{X}$ is first projected by a depth-wise $1{\times}1$ adapter $\boldsymbol{\Phi}\in\mathbb{R}^{C\times d}$, then patchified into $\{\mathbf{x}_{i}\}_{i=1}^{N}$. Thus $\mathbf{Z}^{(0)}$ explicitly concatenates [SINK], class-specific prompts, and embedded patch tokens.
Each spectral prompt $\mathbf{p}_{c,g}$ acts as a “spectral anchor,” interacting only with patches from its assigned band group $\mathcal{B}_{g}$. This design enforces group-aware attention routing and enhances spectral interpretability compared to generic prompts [3, 8].
II-C Sink-Aware Attention Routing
At transformer layer $\ell$ and head $h$, attention is computed as:
@@ II-C Sink-Aware Attention Routing
where $\mathbf{Q},\mathbf{K}\in\mathbb{R}^{(1+G+N)\times d}$, and the [SINK] token index is $k_{\mathrm{sink}}=1$.
We identify a subset of low-informativeness heads $\mathcal{H}_{\text{aux}}$ to serve as noise absorbers. Concretely, we compute attention entropy per head and select the bottom $30\%$ in each layer as $\mathcal{H}_{\text{aux}}$. We then apply an attraction loss:
@@
where $\lambda$ is a small scalar (e.g., $10^{-3}$).
We further define the [SINK] attention mass at layer $\ell$, head $h$:
$\mathrm{mass}^{\ell,h}=\tfrac{1}{N}\sum_{i=1}^{N}\mathbf{A}^{\ell,h}_{i,k_{\mathrm{sink}}},$ (4)
with Sink-Rate as the mean mass across layers/heads, and Sink-Consistency as the average Spearman correlation of saliency distributions across layers. These metrics quantitatively measure collapse mitigation.
II-D 3-D Prompt-CAM Generation
After the final layer $L$, prompt-to-patch attention from discriminative heads $\mathcal{H}_{\star}=\{1,\dots,H\}\setminus\mathcal{H}_{\text{aux}}$ is aggregated:
@@
where $g(\lambda)$ is the spectral group index of band $\lambda$. This yields a spatial–spectral cube $\mathbf{T}\in\mathbb{R}^{H\times W\times C}$.
We derive a spatial heatmap and spectral curve as:
@@
Both are obtained in a single forward pass, without gradients.
II-E Training Objective and Complexity
The total objective combines task-specific loss $\mathcal{L}_{\text{task}}$ (cross-entropy or MSE) and $\mathcal{L}_{\text{sink}}$:
@@ II-E Training Objective and Complexity
During training, only (i) spectral prompts ($K\times G\times d$), (ii) the [SINK] token ($d$), and (iii) the $1\times 1$ spectral adapter ($C\times d$) are updated, adding fewer than 1% parameters (0.8M vs 86M for ViT-B/16) and $\sim$2.1% FLOPs.
III Experiments
III-A Experimental Setup
@@
III-B Main Results
We compare FOCUS against competitive baselines on two challenging hyperspectral benchmarks: HyperLeaf 2024 and Tomato BLS. The quantitative results are summarized in Table I. For classification, we evaluate performance using Accuracy (Acc), band-level interpretability (BIO@5), spatial IoU, and AUPRC. For regression tasks on HyperLeaf2024, we report Mean Squared Error (MSE) and $R^{2}$. To ensure fair comparison, all methods utilize the same frozen ViT-B/16 (DINOv2) backbone. As shown in the table, FOCUS consistently outperforms baselines across all metrics. It demonstrates not only strong predictive accuracy but also superior interpretability in both spatial and spectral dimensions. On HyperLeaf 2024, our method achieves substantial gains. Specifically, it improves BIO@5 and IoU by over 15% compared to the best transformer-based alternatives. Similarly, on Tomato BLS, FOCUS yields sharper lesion localization and more accurate band attribution. This highlights its robustness against domain shifts caused by disease. The advantages extend to regression tasks as well. FOCUS reduces MSE by approximately 10–15% and increases $R^{2}$ by 5–8 points across all targets. Collectively, these results validate FOCUS as a robust and unified framework for accurate HSI modeling.
TABLE I: Main results across hyperspectral benchmarks. For classification (HyperLeaf2024, Tomato BLS), we report accuracy (Acc), BIO@5, IoU, and AUPRC. For regression (HyperLeaf2024: GrainWeight, $G_{sw}$, $\Phi_{\mathrm{PSII}}$), we report mean squared error (MSE, $\downarrow$) and $R^{2}$ ($\uparrow$). Best results are in bold.
Method HyperLeaf2024 (Cls) Tomato BLS (Cls) GrainWeight $G_{sw}$ $\Phi_{\mathrm{PSII}}$
Acc BIO@5 IoU AUPRC Acc BIO@5 IoU AUPRC MSE $R^{2}$ MSE $R^{2}$ MSE $R^{2}$
Grad-CAM [1][ICCV’17] 82.1 0.33 0.41 0.53 79.3 0.28 0.39 0.51 0.142 0.62 0.158 0.59 0.167 0.57
SANet [11][TIP’20] 83.0 0.36 0.43 0.56 80.5 0.29 0.41 0.52 0.138 0.64 0.152 0.61 0.161 0.60
Prompt-CAM [3][CVPR’25] 84.5 0.38 0.46 0.59 81.2 0.34 0.43 0.58 0.134 0.66 0.148 0.63 0.156 0.62
DiffCAM [12][CVPR’25] 85.3 0.41 0.49 0.62 82.4 0.36 0.47 0.61 0.128 0.69 0.145 0.65 0.150 0.64
FOCUS (Ours) 87.0 0.56 0.58 0.72 84.6 0.47 0.54 0.70 0.115 0.75 0.132 0.72 0.141 0.70
TABLE II: Efficiency and inference overhead. FOCUS adds minimal overhead and requires no gradients.
Method Params (%) FLOPs (%) Grad Latency (ms)